OpenStudy (anonymous):

please help! at a dog park there are dogs and their masters. counting heads there are 20 and counting legs there are 62 .. how many dogs and masters are there?

OpenStudy (mertsj):

heads = dogs + masters legs = 4 times dogs + 2 times masters With letters: h=d+m l=4d+2m Now substitute 20 for h and 62 for l and solve.

OpenStudy (anonymous):

can youwork this out for me please.. i ll get it next time around

OpenStudy (mertsj):

d+m=20 4d+2m=62 From the first equation, m = 20-d Replace m in the second equation with 20-d and solve for d